OpenStudy (anonymous):

How many different outfits can you make with 3 pairs of shoes, 4 pairs of pants, 5 shirts and 2 hats?

OpenStudy (anonymous):

3*4*5*2 = 120

OpenStudy (anonymous):

This is just the application of Rule of product ( http://en.wikipedia.org/wiki/Rule_of_product) of combinatorics. With every pair of shoes we can choose to wear any pants and any shirt and any hat so total number of ways of doing this will be \[3 \times 4 \times 5 \times 2 = 5!=120 \]